# General Flow as Foundation Affordance for Scalable Robot Learning

### v1

@@ B.2 From RGBD Human Videos
w/o robustness augmentation: these variants omit three types of technical augmentations (Scale Rebalance, Hand Mask Augmentation, Query Point Sampling Augmentation) to determine their impact on the model’s prediction accuracy in our benchmark.
Experiment Result: The results of the ablation study are summarized in Table IV. For all variants except those without robustness augmentation, there is a noticeable degradation in model performance. In regards to the ablation of the three technical augmentations, it is evident that they do not detrimentally affect benchmark performance. Notably, the hand mask augmentation even significantly enhances in-domain prediction, which is an interesting observation.
| full | 0.0358 | 0.0477 | ||||
| w/o Text EarlyFusion | 0.0370 | 0.0495 | ||||
| w/o Scale Normalization (relative) | 0.0376 | 0.0504 | ||||
| w/o Scale Normalization (absolute) | 0.0381 | 0.0512 | ||||
| w TDN Scale Normalization | 0.0374 | 0.0500 | ||||
| w SDN Scale Normalization | 0.0377 | 0.0510 | ||||
| w (weight of scale-loss) | 0.0368 | 0.0493 | ||||
| w/o central crop | 0.0399 | 0.0538 | ||||
[…]
| w/o HM Augmentation | 0.0366 | 0.0488 | ||||
| w/o QPS Augmentation | 0.0358 | 0.0477 |
-G Real-World Environment Setting
In our real-world experiment (Figure 9), we select 8 objects, including rigid, articulated, and soft bodies, as featured in our human video resources. We manually define multiple tasks and their corresponding success conditions for each object, resulting in a total of 18 distinct tasks. For a complete listing of these tasks, please refer to the content in the main paper (Section VI-B). For “Refrigerator” and “Drawer”, we refer to them as ”Storage Furniture” in the instructions. “Box” is also referred to as “Safe” since it can be seen as an approximation of an atypical design of “Safe”.
These objects are arranged into 6 scenes, as depicted in Figure 9. For objects that are movable, such as “Mug” and “Toy Car”, we randomly adjust their positions and orientations to add variability. It is worth noting that our experimental setup more accurately mirrors practical real-world scenarios compared to previous studies like Eisner et al. [22]. Our setup features diverse scenes and eliminates the necessity for clean object segmentation (Figure 10). This resemblance underscores the robustness and stability of our system in real-world scenarios.
@@
-H Real-World Baseline
To the best of our knowledge, Bahl et al. [4] is the only open-source work that shares a similar setting with ours, which involves learning a low-level affordance model directly from real-world human videos. We deploy this model in our experimental environment, using ’safe’ and ’drawer’ as text prompts for the grounding module [54]. Figure 11 shows the visualization of the affordance prediction. Although it provides semantically meaningful predictions to some extent, it is limited by: (1) inadequate generalization for accurate motion direction prediction; (2) providing only 2D guidance without depth information; and (3) significant disturbances caused by embodiment occlusions. Given that the predicted post-grasp trajectories do not offer sufficient 3D guidance for closed-loop execution, we refrain from further robot execution trials.
-I Development of the Robotic System
Figure 12 shows a snapshot of our real-world deployment setup. We use a RealSense D455 RGBD camera to capture point cloud streams at a resolution of 1280 720, which is lower than the 1920 1080 resolution used in HOI4D [55]. Consequently, we opt for 0.01cm voxel downsampling during deployment, as opposed to the 0.02cm used in model training. The calibration parameters for our camera-robot system are (qw=0.911, qx=-0.015, qy=0.410, qz=-0.032) for orientation and (x=-0.265, y=0.260, z=1.095) for position. This configuration mimics a human ego-view manipulation perspective, beneficial not only for minimizing inference domain-gap but also aligning with practical applications in mobile robots. We utilize the RealSense camera’s ROS driver for data acquisition.
@@
-J Heuristic Policy Derivation
@@
Table V details the average inference latency of each component in our pipeline, based on 10 measurements of the ”open(grasp) Refrigerator” task. The significant bottleneck is the FastSAM segmentation, which contributes to 85.7% of the latency. This highlights the need for more efficient open-world segmentation models in future work. Without FastSAM segmentation, general flow prediction is not the sole system bottleneck; stream acquisition of point clouds also presents substantial room for improvement.
-L Failure Case Analysis
@@
Failure case videos are available on project website. In conclusion, most of these problems are addressable in future deployments. We systematically summarize these limitations and potential improvements in the next section.
-M Future Directions
We discuss the limitations of our current system and outline potential directions for future improvements:
Scaling Up with More Data: Our model still lacks sufficient geometric guidance for complex tasks. Additionally, current data lacks fine-grained language control, for example, summarizing both revolute and prismatic structures as ”Storage Furniture.” More extensive data training could resolve these problems. Fortunately, the release of large-scale, semantically rich human manipulation datasets like [34] coincides with our work provide a rich resource for future research. Moreover, RGB datasets [16, 33] with depth estimation techniques [10, 50] could also be valuable training resources.
Enhanced Downstream Policy Derivation: Our current system relies on a basic heuristic policy based on general flow prediction. This approach has limitations, such as failing when no points surround the gripper. Future developments could leverage both few-shot imitation learning [85] and reinforcement learning [4] for better policy derivation.
Advancements in Deployment Techniques: There is considerable scope for improvement in real-world deployment. Incorporating wrist cameras and multi-view cameras could help alleviate occlusion issues and enhance point cloud quality. More accurate camera calibration and end-effector control could also be helpful for performance enhancement.
-N Codebases

### v2

@@ F.1 Dataset Experiment Setting
w/o robustness augmentation: these variants omit three types of technical augmentations (Scale Rebalance, Hand Mask Augmentation, query points sampling Augmentation) to determine their impact on the model’s prediction accuracy in our benchmark.
The results of the ablation study are summarized in Table 4. For all variants except those without robustness augmentation, there is a noticeable degradation in model performance. In regards to the ablation of the three technical augmentations, it is evident that they do not detrimentally affect benchmark performance. Notably, the hand mask augmentation even significantly enhances in-domain prediction, which is an interesting observation.
| Test-ADE (w/o hand) | Test-FDE (w/o hand) | |
| Full | 3.58 | 4.77 |
| w/o Text EarlyFusion | 3.70 | 4.95 |
| w/o Scale Normalization (relative) | 3.76 | 5.04 |
| w/o Scale Normalization (absolute) | 3.81 | 5.12 |
| w/ TDN Scale Normalization | 3.74 | 5.00 |
| w/ SDN Scale Normalization | 3.77 | 5.10 |
| w/ (weight of scale-loss) | 3.68 | 4.93 |
| w/o central crop | 3.99 | 5.38 |
| w/o Scale Rebalance | 3.59 | 4.78 |
| w/o HMA Augmentation | 3.66 | 4.88 |
| w/o QPS Augmentation | 3.58 | 4.77 |
F.2 Real-Robot Ablation Study
| open Safe | close Drawer | Avg | |
| full | 9 / 10 | 10 / 10 | 95% |
| w/o Scale Rebalance | 7 / 10 | 8 / 10 | 75% |
| w/o HMA Augmentation | 8 / 10 | 6 / 10 | 70% |
| w/o QPS Augmentation | 5 / 10 | 4 / 10 | 45% |
We also investigate the role of designed augmentations in enhancing the robustness of zero-shot transfer via real-robot ablation study. We select two representative tasks (“open Safe” and “close Drawer”) for evaluation. The ablation comparisons included:
The results in Table 5 indicate that each augmentation significantly contributes to robust zero-shot execution.
Notably, hand augmentation markedly affects tasks with substantial embodiment occupancy and occlusion, such as “close Drawer”. Query point sampling augmentation emerges as particularly influential. Currently, the PointNeXt [20] architecture inherently couples the extraction of query point features, leading to a reliance on query point sampling augmentation in our framework. We anticipate that future advancements in disentanglement architecture within 3D learning will address this issue thoroughly.
Appendix G Real-Robot Experiment
G.1 Real-World Environment Setting
In our real-world experiment (Figure 10), we select 8 objects, including rigid, articulated, and soft bodies, as featured in our human video resources. We manually define multiple tasks and their corresponding success conditions for each object, resulting in a total of 18 distinct tasks. For a complete listing of these tasks, please refer to the content in the main paper (Section VI.B). For “Refrigerator” and “Drawer”, we refer to them as ”Storage Furniture” in the instructions. “Box” is also referred to as “Safe” since it can be seen as an approximation of an atypical design of “Safe”.
These objects are arranged into 6 scenes, as depicted in Figure 10. For objects that are movable, such as “Mug” and “Toy Car”, we randomly adjust their positions and orientations to add variability. It is worth noting that our experimental setup more accurately mirrors practical real-world scenarios compared to previous studies like Eisner et al. [6]. Our setup features diverse scenes and eliminates the necessity for clean object segmentation (left of Figure 11). This resemblance underscores the robustness and stability of our system in real-world scenarios.
@@ G.1 Real-World Environment Setting
G.2 Real-World Baseline
To the best of our knowledge, Bahl et al. [11] is the only open-source work that shares a similar setting with ours, which involves learning a low-level affordance model directly from real-world human videos and could perform zero-shot prediction. We deploy this model in our experimental environment, using ’safe’ and ’drawer’ as text prompts for the grounding module [66]. Right of Figure 11 shows the visualization of the affordance prediction. Although it provides semantically meaningful predictions to some extent, it is limited by: (1) inadequate generalization for accurate motion direction prediction; (2) providing only 2D guidance without depth information; and (3) significant disturbances caused by embodiment occlusions. Given that the predicted post-grasp trajectories do not offer sufficient 3D guidance for closed-loop execution, we refrain from further robot execution trials.
G.3 Development of the Robotic System
The heuristic policy based on close-loop flow prediction is demonstrated in Algorithm 1. Figure 7 shows a snapshot of our real-world deployment setup. We use a RealSense D455 RGBD camera to capture point cloud streams at a resolution of 1280 720, which is lower than the 1920 1080 resolution used in HOI4D [15]. Consequently, we opt for 0.01cm voxel downsampling during deployment, as opposed to the 0.02cm used in model training. The calibration parameters for our camera-robot system are (qw=0.911, qx=-0.015, qy=0.410, qz=-0.032) for orientation and (x=-0.265, y=0.260, z=1.095) for position. This configuration mimics a human ego-view manipulation perspective, beneficial not only for minimizing inference domain-gap but also aligning with practical applications in mobile robots. We utilize the RealSense camera’s ROS driver for data acquisition.
@@ G.3 Development of the Robotic System
G.4 Heuristic Policy Derivation
@@ G.5 Inference Latency
Table 6 details the average inference latency of each component in our pipeline, based on 10 measurements of the ”open(grasp) Refrigerator” task. The significant bottleneck is the FastSAM segmentation, which contributes to 85.7% of the latency. This highlights the need for more efficient open-world segmentation models in future work. Without FastSAM segmentation, general flow prediction is not the sole system bottleneck; stream acquisition of point clouds also presents substantial room for improvement.
G.6 Failure Case Analysis
@@
Failure case videos are available in https://general-flow.github.io. In conclusion, most of these problems are addressable in future deployments. We systematically summarize these limitations and potential improvements in the next section.
Appendix H Related Works of Real-World General Robot Learning
Research in general-purpose robotic manipulation in real-world settings is constantly evolving, with a focus on integrating Large Language Models (LLMs) for high-level planning [68, 69, 70, 71, 72, 73] and exploring direct actionable guidance through LLMs [74], albeit with challenges due to overlooked physics dynamics. The development of large models for direct low-level control [75, 76, 77, 78] faces scalability issues due to intensive data requirements [3, 4]. This highlights the need for a training framework that achieves a balance between actionable output and scalability. In this work, we achieve this through a scalable robot learning approach based on general flow prediction.
Appendix I Limitations
The diversity and volume of our data are currently limited, which restricts our model’s ability to provide adequate guidance for complex tasks. A future direction could involve utilizing larger RGBD datasets [79] or RGB datasets [35, 80] combined with depth estimation techniques [81, 32] to address these issues.
We currently employ a Conditional VAE to capture the multimodal distribution of human behavior in our dataset. However, this architecture may lack sufficient expressive power when dealing with more challenging scenarios, such as larger and more diverse datasets. To address this limitation, the Diffusion model [82] could be a potential solution.
In addition, the manual grasping process could be replaced with heuristic policies such as those described in AVDC[32], or through more advanced and automated methods like DexFunc[14], MoKa[13], and ManipVQA[83] in the future. It is also feasible to expand the representation from object-centric affordance to universal keypoints motion, similar to what is described in ATM[62].
The initially adopted ICP-based heuristic policy in a zero-shot manner might limit the completion of tasks, particularly those requiring contact-rich manipulation. Future studies could fine-tune policies based on general flow predictions within a few-shot imitation learning setting, such as Track2Act[63], or use general flow as an action constraint [29] for sample-efficient reinforcement learning, akin to VRB [11].
Appendix J Codebases